solve by system of addition
3x - 2.5y =7.125
2.5x - 3Y =7.3125

Answer by checkley77(12844)

3x - 2.5y =7.125 multipl by -2.5
2.5x - 3Y =7.3125 multiply by 3 & add
-7.5x+6.25y=-17.8125
7.5x-9y=21.9375
------------------------
0x-2.75y=4.125
y=4.125/-2.75
y=-1.5 ans.
3x-2.5*-1.5=7.125
3x+3.75=7.125
3x=7.125-3.75
3x=3.375
x=3.375/3
x=1.125 ans.
Proof:
2.5*1.125-3*-1.5=7.3125
2.8125+4.5=7.3125
7.3125=7.3125